# ArxEval: Evaluating the Retrieval and Generation of Language Models for Scientific Papers

### v1

@@ 1 Introduction
In this paper, we present a comprehensive study evaluating the extent of hallucination in LLMs under domain-specific prompting, with a particular focus on scientific literature. We develop and implement a systematic evaluation pipeline to assess fifteen prominent open-source LLMs: Qwen 2.5 [9], Gemma 2 [10], Llama 3 [11], Phi 3 [12], Orca 2 [13], Mistral v-0.3 [[14], Deepseek-llm [15], Olmo-2 [16], Mistral-Nemo [17], Eurus-2 [18], and Solar-Pro [19]. Our evaluation utilizes the ArXiv dataset [20] as the primary source of scientific articles, providing a robust foundation for assessing model performance in academic contexts.
@@ 2.5 Hallucinations in Academic Reference Generation
Academic reference generation represents a critical challenge, with studies demonstrating that even state-of-the-art models frequently generate fabricated or inaccurate citations [31]. This limitation underscores the urgent need for continued research in hallucination mitigation, particularly in tasks where factual accuracy is paramount. In addressing these challenges, our work specifically focuses on Domain-Specific Biases by utilizing over 150 categories of papers from the ArXiv repository, providing a comprehensive evaluation across diverse academic domains.
@@ 3 Dataset
Figure 1: Number of titles from each subject.
@@
The Jumbled Title task includes 528 jumbled titles, distributed evenly among the 176 categories of the ArXiv dataset, Table 1 presents a few examples of the dataset used for the Jumbled Title task.
@@ 5 Results
Figure 5: DOIs Found for each model vs DOIs not for each model

### v2

@@ 1 Introduction
In this paper, we present a comprehensive study evaluating the extent of hallucination in LLMs under domain-specific prompting, with a particular focus on scientific literature. We develop and implement a systematic evaluation pipeline to assess fifteen prominent open-source LLMs: Qwen 2.5 [9], Gemma 2 [10], Llama 3 [11], Phi 3 [12], Orca 2 [13], Mistral v-0.3 [14], Deepseek-llm [15], Olmo-2 [16], Mistral-Nemo [17], Eurus-2 [18], and Solar-Pro [19]. Our evaluation utilizes the ArXiv dataset [20] as the primary source of scientific articles, providing a robust foundation for assessing model performance in academic contexts.
@@ 2.5 Hallucinations in Academic Reference Generation
Academic reference generation represents a critical challenge, with studies demonstrating that even state-of-the-art models frequently generate fabricated or inaccurate citations [31]. This limitation underscores the urgent need for continued research in hallucination mitigation, particularly in tasks where factual accuracy is paramount. In addressing these challenges, our work specifically focuses on Domain-Specific Biases by leveraging over 150 categories of papers from the ArXiv repository, providing a comprehensive evaluation across diverse academic domains.
Figure 1: Number of titles from each subject.
@@ 3 Dataset
The Jumbled Title task includes 528 jumbled titles, distributed among the 176 categories of the ArXiv dataset, Table 1 presents a few examples of the dataset used for the Jumbled Title task.
@@ 5 Results
Figure 5: DOIs Found for each model vs DOIs not for each model